a populatpion of 1,000 students has an average weekly allowance of =350 and standard deviation of =56.13 .what is the probability that a random sample of size =30 will have an average weekly allowance between 335 and 360?



1
Expert's answer
2022-03-31T02:46:00-0400

We have a normal distribution, "\mu=350, \sigma=56.13."

Let's convert it to the standard normal distribution, "z=\cfrac{\bar{x}-\mu}{\sigma/\sqrt{n}};"

"z_1=\cfrac{335-350}{56.13/\sqrt{30}}=-1.46, \\z_2=\cfrac{360-350}{56.13/\sqrt{30}}=0.98,"


"P(335<X<360)=P(-1.46<Z<0.98)=\\=P(Z<0.98)-P(Z<-1.46)="

"=0.83646-0.07215=0.76431" ​(from z-table).